OpenStudy (anonymous):

simplify (x^1/2)^1/6

OpenStudy (anonymous):

X^1/12

OpenStudy (anonymous):

\[(x ^{a})^{b} = x ^{a \times b}\]

OpenStudy (anonymous):

imagine that you'd have this : \[(x^2)^3\] What that means is actually this, right? \[x^2 * x^2 * x^2 = x*x*x*x*x*x = x^6\] and so you find that x^6 = (x^2)^3 which proves the rule